OpenStudy (leozap1):

Fan & Medal. Jeffrey earned a gross income of $51,280 last year. He made $1,746.83 in student loan interest deductions, donated $2,914 to his favorite charities and paid $3,941.47 in home mortgage interest. Jeffrey claims a standard deduction of $11,400 for himself and his non-working spouse. If their exemption is $7,300, what is their taxable income? A. $26,891.70 B. $23,977.70 C. $35,377.70 D. $30.833.17

OpenStudy (anonymous):

what have u done so far?

OpenStudy (leozap1):

I don't know what to do.

OpenStudy (anonymous):

I'm re-reading this

OpenStudy (leozap1):

Do I subtract the amounts listed? I did that and got B.

OpenStudy (leozap1):

$51,280-$1,746.83-$2,914-$3,941.47 -$11,400-$7,300

OpenStudy (anonymous):

srry! I got logged out

OpenStudy (leozap1):

Okay, no problem lol. Is that how I solve this problem though? It seems too easy lol.

OpenStudy (anonymous):

hmh. let me see

OpenStudy (anonymous):

were's the answer you got?

OpenStudy (leozap1):

I get $23,977.70.

Directrix (directrix):

Jeffrey claims a standard deduction of $11,400 for himself and his non-working spouse. He took the standard deduction and did not itemize. His loan interest and charity contributions will not be subtracted from his gross income because of that.

OpenStudy (leozap1):

So that would give me 28,638.53 but that's not an answer?

Directrix (directrix):

51,280 - 11,400 - 7,300 = is what I think for his adjusted gross income.

OpenStudy (anonymous):

yes,but you must do another step

OpenStudy (anonymous):

which will give what?

OpenStudy (leozap1):

That's 32,580. My answer choices are A. $26,891.70 B. $23,977.70 C. $35,377.70 D. $30.833.17

Directrix (directrix):

I don't know what that is not correct. I need to look on the IRS site to see if any "special" costs can also be deducted from the gross income when the standard deduction is taken.

Directrix (directrix):

Does anybody know if student loan interest is directly subtracted from the gross income even if a person takes the standard deduction and does not itemize. If so, 51,280 - 11,400 - 7,300 -1746= very close to D. $30.833.17 That may just be happenstance so don't view it as an answer.

Directrix (directrix):

"You might be paying back loans you took to finance higher education. If so, you could qualify to deduct up to $2,500 of interest per return per year. You can claim the student-loan interest deduction as an adjustment to income. You don't need to itemize deductions to claim it." From H&R Block website I think that the student loan interest is subtracted from the gross income even though Jeff took the standard deduction.

Directrix (directrix):

I am sticking with this as my best idea of an answer based on the above information. 51,280 - 11,400 - 7,300 -1746= very close to D. $30.833.17 @leozap1

Directrix (directrix):

I forgot to subtract the $.83 of the loan. Now an answer match is possible. 51,280 - 11,400 - 7,300 -1746.83= $30.833.17

OpenStudy (leozap1):

Okay, thank you! I will let you know if it was the right answer is when I find out. :)

OpenStudy (leozap1):

D was the correct answer, thank you! :D

Directrix (directrix):

You are welcome. I learned about student loan deductions from reading about this problem.